OpenStudy (anonymous):

Which equation can be used to find the answer? After two months, Rachel has saved enough money to buy an MP3 player. During the second month, she saved $87 and she pays $199 for the MP3 player. How much did Rachel save during the first month? A. m = 87 • 2 m = 174 $174 B. 199 + 87 = m m = 286 $286 C. m + 87 = 199 m = 112 $112 D. m – 87 = 199 m = 287 $287 PLease help and i do give medals

OpenStudy (anonymous):

@96Neko

OpenStudy (anonymous):

C. m + 87 = 199 m = 112 $112 You have to do the inverse operation, or subtraction: 199 - 87 = 112

OpenStudy (anonymous):

Ok Thanks.. Will you help me with more?

OpenStudy (anonymous):

I just noticed my explanation is confusing so let me rephrase it... the inverse operation is subtraction, so you subtract 87 from 199 And yeah, I'll help you with more

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

Which equation can be used to find the answer? Christina climbs to the top of Mount Graham in Arizona. On the first day, she climbs to an elevation of 1029 ft. The elevation of Mount Graham is 10,724 ft. How many total feet of Mount Graham’s elevation does Christina climb on the other days? A. e + 1029 = 10,724 e = 9695 9695 ft B. e = 1029 · 2 e = 2058 2058 ft C. e = 10,724 ÷ 2 e = 5362 5362 ft D. e – 1029 = 10,724 e = 11,753 11,753 ft

OpenStudy (anonymous):

Idk why there are them little thingys are there

OpenStudy (anonymous):

I can't read that, can you try to take a screenshot of the problem ad attach it as a file

OpenStudy (anonymous):

Ok. Hang on just second

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

A Christina climbed 1029 ft. of the mountain on the first day, so if you subtract it from the mountain's total height, you'll get what she climbed in total on the other days.

OpenStudy (anonymous):

Ok thanks.. Let me get you a medal i have 3 more

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

D This problem is basically a subtraction problem, because it's subtracting 96 from whatever the number was originally. The inverse of subtraction is addition, so: 96 + 172 = 268 is your answer

OpenStudy (anonymous):

Thanks!:)

OpenStudy (anonymous):

No problem. Do you have any more?

OpenStudy (anonymous):

Yeah 1 more after this 1 i'm fixing to post

OpenStudy (anonymous):

C This is an addition problem, because you're adding x number of stamps to completely fill the book. So we would do subtraction to find out what x is. 54 - 38 = 16

OpenStudy (anonymous):

Thank you again

OpenStudy (anonymous):

D This is subtraction, because 23 were sold. So you have to add 23 to the amount she has left: 112 + 23 = 135

OpenStudy (anonymous):

Thank you! i got a 100 Thanks

OpenStudy (anonymous):

No problem :)